Use technology to find the​ P-value for the hypothesis test described below. The claim is that for a smartphone​ carrier's data speeds at​ airports, the mean is μ=17.00 Mbps. The sample size is n=29 and the test statistic is t=2.693 ​what is the p-value

​(Round to three decimal places as​ needed.)
Answer #1

Solution,

Given that,

n = 29

degrees of freedom = n - 1 = 29 - 1 = 28

Test statistic = t = 2.693

P(t > 2.693)

= 1-P (t < 2.693)

= 1 - 0.994

P-value = 0.006